

********************************************************************
UARS WEEKLY STATUS REPORT - 26 March 2004
********************************************************************
***********
SOLSTICE
***********
This week, SOLSTICE continued to operate in good health gathering solar
data on all available orbital tracking periods. On average, we received
62% of each day's data amounting to about 588 minutes of solar
observing time per day. This noticable increase in data capture was
due to the implementation of 3-TDRSS operations.
***********
SUSIM
***********
SUSIM successfully gathered its daily mid (1.1 nm) resolution, Mg II,
occultation, solar pointing offset and high resolution continuous
monitoring scans for periods of UARS real-time telemetry or onboard
tape recording.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 13 March 2004 (Orbit 68413, GMT Day 073) through
19 March 2004 (Orbit 68517, GMT Day 079).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 075/2133 68456 FULLTIME TO DAY-ONLY OPS
PEM AXIS 075/2133 68456 POWER OFF
HALOE 076/1411 68466 POWER ON
HALOE 078/1449 68497 POWER OFF (HI TEMPS)
HALOE 078/1630 68498 POWER ON
The beta angle ranged between the angles of 59.4 degrees to 40.7
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 29 March 2004 (DOY 089).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.27 to 1.29 degrees C. Additional
battery performance data for each battery on 12 March and 19 March
2004 are provided in Appendix D.
The clock error ranged from 10.4 to -10.4 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 01
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 04
PEM = 02
HALOE = 03
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 04
Two (2) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Four (4) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT transitioned the CMS to the new software on 15 March 2004.
All loads are being built on this new version without incident. This
was done to prepare for the start of 3-TDRS operations, which began on
17 March 2004. The FOT is now taking 35-38 events per day (up from 26-
28 per day under 2-TDRS operations) and now collecting data for an
additional 18-20% more of each orbit. This 3-TDRS mode will be the
nominal mode from now on. This new mode will be closely examined to
see if any slight modifications may be needed in the future, but none
are planned at this time. The FOT would like to once again thank all
those involved in making this happen and thank science for being
patient with this effort. Any feedback from science on how this mode
is working for them would be greatly appreciated.
MEETINGS
The FOT held a telecon with WSC personnel on 18 March 2004 to discuss
the capability of automatically sending late acquisition GCMRs. This
capability will be implemented starting 1600Z on 22 March 2004.
The FOT also attended a meeting on 19 March 2004 with the HRDI team to
discuss the attitude error seen in their data. Several data sets will
be reviewed to determine the cause of this error and how it can be
corrected.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 29 March 2004 (DOY 089). This
maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 19 March 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
68417 073/0805 04-080 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
68450 075/1122 04-081 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
68456 075/2133 04-082 PEM FULLTIME TO DAY-ONLY OPS
68456 075/2133 04-083 PEM AXIS POWER OFF
68466 076/1411 04-084 HALOE POWER ON
68493 078/0832 04-085 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
68497 078/1449 04-086 HALOE POWER OFF (HI TEMPS)
68498 078/1630 04-087 HALOE POWER ON
68511 079/1401 04-085 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 19 March 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
HATBAT Y-HI BOLOMETER AREA TEMP 68494
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 19 March 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
68417 073/0805 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
68450 075/1122 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
68493 078/0834 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
68511 079/1401 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 19 March 2004
12 March 2004 - GMT Day 072 Beta = 60.2 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.68/-4.94 +00.0/+00.0 07.0 0.00 0.0
2 +11.2/-022.4 +3.87/+2.60 +22.0/-15.2 26.1 1.02 17.8
3 +11.2/-033.6 +2.91/+1.65 +21.2/-15.2 25.9 1.02 17.5
19 March 2004 - GMT Day 079 Beta = 40.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.31 +00.0/+00.0 07.0 0.00 0.0
2 +11.2/-095.2 +5.79/+4.50 +27.2/-15.2 24.8 1.04 21.5
3 +22.4/-078.4 +4.82/+3.55 +26.4/-15.2 24.6 1.04 21.0
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None This Report Period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 19 March 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-015 FOT 68430 074/0401 35722 Late Acquisition
04-014 FOT 68470 076/1953 30832 Generic Late Acquisition
04-014 FOT 68504 079/0245 30832 Generic Late Acquisition
04-014 FOT 68506 076/0521 30832 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-014 FOT See List Mar 04 30832 Generic Late Acquisition
04-015 FOT 68430 074/0401 35722 Late Acquisition